# SoK: Consensus for Fair Message Ordering

### v1

@@ Abstract
Distributed ledger systems, such as blockchains, rely on consensus protocols that constantly commit messages in an agreed order for processing. In practice, message ordering within these systems is often reward-driven. This raises concerns about fairness, particularly in decentralized finance applications, where nodes can exploit transaction orders to maximize rewards (Maximal Extractable Value, MEV). This paper provides a structured review of consensus protocols that order messages with different approaches, especially focusing on the ones that promote order fairness, using methods including First-In-First-Out (FIFO), random, and blind ordering. We review the challenges and trade-offs of deriving fair message ordering in a Byzantine fault-tolerant setting, and summarize the key steps for making a fair message ordering consensus protocol. We introduce a design guideline, with which we propose a performance optimization to the state-of-the-art FIFO ordering protocol Themis [1]. This work establishes a unified framework for accessing and enhancing fairness in distributed ledger systems.
I Introduction
@@
Consensus protocols in distributed ledger systems establish agreements on the messages (or transactions, used interchangeably in this paper) to process in some specific orders. Decentralized applications, particularly those utilizing smart contracts, often result in different states if messages are processed in different orders. Since nodes in distributed ledger have significant freedom to censor the messages and decide on an order, they can manipulate the states of decentralized applications by altering the order of message processing. In practice, nodes frequently manipulate message ordering to maximize rewards. For example, in major blockchains like Bitcoin [2] and Ethereum [3], block proposers typically favor transactions with higher fees. Consequently, transactions with lower fees experience longer delays, leading to fairness concerns [4].
@@
More importantly, in decentralized finance (DeFi) applications that rely on smart contracts, nodes have economic incentives to manipulate the order of message processing to generate additional profits known as Maximal Extractable Value (MEV), often resulting in substantial financial losses for users. For example, a node can create a transaction to buy an asset and prioritize it if the node observes a transaction in the message pool that can raise the asset price once processed. As a result of such profiting strategies through transaction ordering manipulation, around 200 million USD are extracted from DeFi users every year [5].
@@
The core challenge in deriving a fair ordering of messages in distributed ledger systems is the inherent lack of synchronization across nodes. As shown in Figure 1, in decentralized networks, each node can independently receive messages from external users, resulting in differing local views of the message pool. Even if all nodes eventually receive the same set of messages by constantly exchanging their message pool, they are likely to receive them at varying times and in differing orders due to network delay and asynchrony [6], making it difficult for the nodes to reach an agreement of a fair message processing order. An easy way to prevent asynchronized message pools is to have a centralized entity receiving messages [7], but it violates the nature of decentralization. Moreover, most distributed ledger systems require Byzantine fault tolerance (BFT), which adds additional considerations to the robustness of the fair ordering rule in designing a fair ordering protocol.
Figure 1: A fair ordering consensus protocol achieves fair order agreements among nodes
@@
Awareness of the message ordering problem has arisen in the past five years and has shown a growing interest as the number of related works arises. While there are proposals for fair ordering consensus under network asynchrony and Byzantine-fault nodes, a comprehensive understanding of the challenges, design choices of these systems is yet to be developed. Therefore, we present the first systematization of knowledge of fair ordering consensus protocols by examining existing fair ordering consensus protocols. The key contributions of our work are as follows:
@@
    We review, to the best of our knowledge, all proposals for message ordering and offer a unified perspective on the requirements and limitations of using these ordering rules.
@@
    A case study of the state-of-the-art FIFO ordering protocol, Themis [1]. Through the fair consensus factory, we are able to optimize Themis to reduce its latency.
@@
The remainder of this paper is structured as follows: Section 2 provides background information on current message ordering approaches and MEV issues in distributed ledgers. Section 3 presents a unified perspective on existing message ordering rules. Section 4 presents the fair consensus factory that adds fairness to consensus protocols. Section 5 provides a case study of Themis regarding its consensus design and a latency optimization provided by the fair consensus factory framework. Section 6 discusses the open challenges in fair ordering consensus for future studies. Finally, Section 7 concludes the paper.
II Preliminaries: Status of Real-world Distributed Ledger Message Ordering
II-A Problem Modeling
Figure 2: Message flow in distributed ledgers
In a distributed ledger, messages are constantly sent from external users to one or more nodes in the system, or generated by nodes within the system. Each successful round of consensus commits a new, ordered set of messages to process at the application layer, as shown in Figure 2. For simplicity, we assume all nodes have equal likelihoods of receiving messages and each node has an equal opportunity to participate in message ordering. That is, messages are neither advantaged nor disadvantaged based on the nodes that receive them.
Nodes are assumed to be rational, meaning they act in their own interests if they detect profitable opportunities from manipulating message ordering. For instance, if a node is selected as the leader in a consensus round, it might manipulate message order for financial gain. Additionally, to align with the common security setting of most existing permissionless distributed ledgers, the system must tolerate up to $f$ Byzantine fault nodes out of $n$ total nodes, where $n\geq 3f+1$. Byzantine nodes may deviate arbitrarily from the protocol without preventing honest nodes from reaching consensus. Broadly speaking, this message-ordering challenge applies to any general models of Byzantine fault-tolerant state machine replication [8] systems that require consensus on message ordering.
These systems can use different data structures for state representation, widely used data structures include a chain of message blocks (i.e. blockchain), a directed acyclic graph (DAG) of individual messages, or a DAG of message blocks.
II-B Fairness in Blockchain and DAG-based Ledger
Blockchain and DAG are two primary data structures for storing messages in distributed ledgers. In blockchains, nodes collectively commit one block at a time per consensus round. Given the limited throughput of blockchains, demanding blockchains with high input message rates often results in high latency of committing messages. Without fair ordering mechanisms, transactions with lower fees usually experience longer waiting times, as high-fee transactions are prioritized.
DAG-based systems, by contrast, support greater concurrency, offering higher throughput and lower latency [9]. Although these systems do not enforce specific ordering rules, in reality, they are generally more likely to commit messages in the order they are received, resulting in a FIFO-like effect. Therefore, DAG-based systems are fairer than blockchains in the sense that the latency to commit messages is lower than blockchains due to high throughput, even though they pay low transaction fees. However, nodes can still manipulate message order to exploit MEV opportunities without an enforced fair ordering rule.
II-C Consensus and Message Ordering Implementation in Existing Systems
Distributed ledgers usually rely on Byzantine fault-tolerant (BFT) committee-based consensus mechanisms that impose a total order on messages for sequential execution at the application layer. Li et al. [10] listed the transaction sequencing rules in the consensus protocol implementation of a number of mainstream distributed ledger systems. It is shown that their consensus implementations use different parameters to order messages, including transaction fees, ages, and locality (prioritize local messages). However, these rules are not enforced, that is, each node can modify these implementations to use their own ordering rules. For example, Avalance [11] is a DAG-based system with a standard implementation that prioritizes older transactions (FIFO). However, rational nodes can run an alternative ordering rule if they see MEV profit opportunities, and the resulting order is still approved by the peers.
II-D Maximal Extractable Value
Decentralized finance is a popular category of application that is built on top of distributed ledgers. Maximal Extractable Value (MEV) is the profit opportunities nodes can get by observing uncommitted transactions in decentralized finance applications [12, 13, 5]. Typically, nodes can front-run and/or back-run transactions by manipulating the message ordering mechanism to place a message in front of or behind certain incoming transactions that, once committed, will change the price of certain cryptocurrencies to gain profit from the price differences [13], or simply imitate the transactions with profit potentials and prioritize the imitated transaction [14].
While some claim MEV has a positive effect on a system [15], the majority believe they are ”attacks” and should be eliminated [12]. Although there are studies to avoid MEV indirectly at the application level [16, 17], the root cause of MEV is that message ordering can be manipulated. With enforced message ordering rules, fair ordering consensus has been identified as an effective solution to the MEV problem [18, 19].
III Message Ordering Rules
@@
This section formalizes different message ordering rules in the Byzantine fault tolerant setting and discusses in detail the requirement of executing these ordering rules and their limitations.
@@
We define a message ordering rule as a function that sorts an unordered set of messages into a specific sequence based on certain metadata. Let $\mathcal{S}^{m}$ and $\mathcal{O}^{m}$ be the family of the sets of $m$ unordered messages and ordered messages ($|S|=|O|=m$), respectively. Let $\mathcal{D}$ denote the set of metadata used to order messages (varies depending on the actual ordering rule) and $D\in\mathcal{D}$. Now we define a message ordering rule function:
@@
That is, a message ordering rule function orders a set of messages $S$, which comes from the aggregation of all local message pools of the network nodes, using some metadata $D$ and outputs the messages in an ordered list ($s.t.\ \forall\ tx\in O,\ tx\in S$). Below, we discuss in detail the types of ordering rules $r$ and the corresponding metadata $D$ used to decide the ordering of the messages in these rules.
III-A FIFO Ordering Rule
@@
However, since messages can be received by different nodes at different times and orders at the network layer, the major challenge of FIFO ordering rules is to derive a fair global message order using these different local views. Moreover, in the Byzantine fault tolerant setting, the rules have to consider that up to $f$ local views are missing or indistinguishably malicious. Two main methods to achieve FIFO ordering under these conditions are ranked voting and median timestamping.
III-A1 FIFO via Ranked Voting
The process of collectively agreeing on an order when each voter node has a different local order preference is, in one simple term, ranked-choice voting. Therefore, there are some proposals to derive the order of messages by collecting local message arrival orders from each individual node and aggregating them [20, 22, 23, 24, 1, 25, 26, 39].
Given an unordered message set $S$, let $O^{S}_{i}$ be the ordered set of messages based on the arrival order of messages locally observed by node $i$. FIFO order via voting is defined as the following ordering rule:
@@ III-A1 FIFO via Ranked Voting
In FIFO via ranked voting, each node’s local order of messages acts as a “ballot” that ranks the messages, and the final ordering is decided by majority preference among nodes. However, this method may not always yield a strict FIFO order due to Condorcet cycles [40], where the relative order of messages creates cyclic preferences among nodes [20]. Figure 3 is a demonstration of how a Condorcet cycle is formed in ranked voting. In this example, the majority of the local views suggest message 1 comes before message 2, message 2 comes before message 3, but also believe message 3 comes before message 1. Even worse, such cycles can be infinitely extended. This raises an attack opportunity of deliberately creating or extending Condorcet cycles in the case where malicious external message senders submit messages to different nodes at carefully designed times, under the condition that nodes in the network do not disseminate local messages timely [27].
@@
Figure 3: An example of Condorcet cycle from ranked voting
@@
To tolerate Condorcet cycles, different relaxed notions of order fairness are proposed. Block order fairness [20] (later referred to as batch order fairness) and similarly differential order fairness [23] are proposed that allow multiple messages to be deemed as received at the same time.
@@
As Condorcet cycles can be arbitrarily long, making sure messages in Condorcet cycles output together also impacts the liveness of the protocol [20]. A solution to this problem is to assume an external synchrony time bound [20], so that when the time of receiving two messages is over the time bound, they must not be in a Condorcet cycle. The drawback of this method is the high latency to confirm a message and the requirement of the additional external synchrony assumption. To improve on that, [27, 28] proposed to use ranked pairs [41] to break Condorcet cycles without violating FIFO order fairness. Themis [1] uses a method called batch-unspooling to commit messages without requiring to observe full cycles.
@@
To summarize, this line of research attempts to handle Condorcet cycles in a way that minimizes their effect on latency. However, taking a step back, there lack of a view of the formation of Condorcet cycles. An attacker can create Condorcet cycles only if nodes do not synchronize the message they receive with their peers timely [27]. Therefore, a potential research direction that can advance FIFO via voting proposals is to reduce the probability of forming Condorcet cycles by requiring more frequent message propagation.
III-A2 FIFO via Timestamping

### v3

@@ Abstract
Distributed ledger systems, such as blockchains, rely on consensus protocols that commit ordered messages for processing. In practice, message ordering within these systems is often reward-driven. This raises concerns about fairness, particularly in decentralized finance applications, where nodes can exploit transaction orders to maximize rewards referred to as Maximal Extractable Value. This paper provides a systematic understanding of consensus protocols that order messages with different approaches, especially focusing on the ones that promote order fairness, using methods including First-In-First-Out (FIFO), random, and blind ordering. We review the challenges and trade-offs of deriving fair message ordering in a Byzantine fault-tolerant setting, and summarize the requirements for making a fair message ordering consensus protocol. We introduce a design guideline, with which we propose a latency optimization to the state-of-the-art FIFO ordering protocol of Themis [1]. This work provides a systematic way for assessing and enhancing message order fairness in blockchain systems.
Index Terms:
blockchain, consensus, ordering, fairness
1 Introduction
@@
Consensus protocols in blockchain systems establish agreements on the messages (or transactions, used interchangeably in this paper) to process in some specific order. Decentralized applications, particularly those utilizing smart contracts, often result in different states if messages are processed in different orders. Since nodes in blockchains have significant freedom to censor the messages and decide on an order, they can manipulate the states of decentralized applications by altering the order of message processing. In practice, nodes frequently manipulate message ordering to maximize rewards. For example, in major blockchains such as Bitcoin [2] and Ethereum [3], block proposers typically favor transactions with higher fees. Consequently, transactions with lower fees experience longer delays, leading to fairness concerns [4].
@@
More importantly, in decentralized finance (DeFi) applications that rely on smart contracts, nodes have economic incentives to manipulate the order of message processing to generate additional profits known as Maximal Extractable Value (MEV), often resulting in substantial financial losses for users [5]. For example, a node can create a transaction to buy an asset and prioritize it if the node observes a transaction in the message pool that can raise the asset price once processed. As a result of such profiting strategies through transaction ordering manipulation, around 200 million USD are extracted from DeFi users every year [5].
Although there are multiple definitions and approaches that label message order fairness in existing literature, a commonality among these definitions is that they are designed to prevent MEV in blockchains, where a single node in the system can decide message ordering out of its own interest.
@@
The core challenge in deriving a fair ordering of messages in distributed ledger systems is the state asynchronization across nodes. As shown in Figure 1, in decentralized networks, each node can independently receive messages from external users, resulting in differing local views of the message pool. Even if all nodes eventually receive the same set of messages by constantly exchanging their message pool, they are likely to receive them at varying times and in differing orders due to network delay and asynchrony [6], making it difficult for the nodes to reach an agreement of a fair message processing order. An easy way to prevent asynchronous message pools is to have a centralized entity receiving messages [7], but it violates the nature of decentralization. Moreover, most distributed ledger systems require Byzantine fault tolerance (BFT), which adds additional considerations to the robustness of the fair ordering rule in designing a fair ordering protocol.
Figure 1: An example consensus protocol that achieves FIFO order agreements among the four nodes. The nodes agree to commit messages in the order of message 1, 2, 3, 4, and 5. The protocol does not guarantee to commit all messages in one agreement. In this example, message 6 is uncommitted in this round of agreement; it stays in the local message pool to get committed in upcoming rounds of agreement.
@@
Awareness of the message ordering problem has arisen in the past five years and has shown a growing interest as the number of related works has grown. While there are proposals for fair ordering consensus under network asynchrony and Byzantine-fault nodes, a comprehensive understanding of the challenges, and design choices of these systems is yet to be developed. Therefore, we present the first systematization of knowledge of fair ordering consensus protocols by examining existing fair ordering consensus protocols. The key contributions of our work are as follows:
@@
    We review proposals for message ordering and offer a unified perspective on the requirements and limitations of using these ordering rules.
@@
    A case study of the state-of-the-art FIFO ordering protocol, Themis [1]. Through the Fair Consensus Factory, we propose an alternative design of Themis that reduces latency.
@@
The remaining paper is structured as follows: Section 2 provides background information on current message ordering approaches and MEV issues in distributed ledgers. Section 3 presents a unified perspective on existing message ordering rules. Section 4 presents the Fair Consensus Factory that adds fairness to consensus protocols. Section 5 provides a case study of Themis regarding its consensus design and a latency optimization provided by the fair consensus factory framework. Section 6 discusses the open challenges in fair ordering consensus for future studies. Finally, Section 7 concludes the paper.
2 Background
2.1 Problem Setting
Figure 2: Message flow in distributed ledgers. Users send messages to nodes, which then commit an ordered batch through consensus.
Distributed ledgers rely on a set of nodes to receive, relay, and agree on an ordered sequence of messages for execution at the application layer (Figure 2). Messages may originate from external clients or be generated internally by nodes. In each consensus round, a new ordered batch of messages is committed. This order determines the state transitions and application-level behavior.
We consider a setting in which nodes have equal opportunity to receive messages and participate in the ordering process. That is, the protocol assumes no geographic, topological, or infrastructural advantage among nodes. However, nodes are rational and may deviate from protocol behavior if doing so increases their individual utility, especially when granted special roles such as the consensus leader.
We adopt the standard Byzantine fault-tolerant (BFT) model. Among $n=3f+1$ participating nodes, up to $f$ may behave arbitrarily, including collusion and deviation from the protocol. Byzantine nodes may deviate arbitrarily from the protocol without preventing honest nodes from reaching consensus. Broadly speaking, this message-ordering challenge applies to any general model of Byzantine fault-tolerant state machine replication [8] systems that require consensus on message ordering.
2.2 Message Ordering in Existing Distributed Ledgers
In most blockchains and DAG-based systems, consensus protocols must determine not just which messages to include but also in what order to process them. Traditionally, Byzantine consensus protocols guarantee agreement on total order, but do not constrain how that order is chosen.
Li et al. [9] surveyed ordering policies in existing distributed ledger systems and found significant variation. Some systems prioritize by transaction fee, others by local reception time or message origin. However, these policies are implementation-level defaults rather than enforced rules. Individual nodes can and often do modify their ordering behavior in pursuit of profit.
For example, Avalanche [10], a DAG-based system, adopts a FIFO policy by default. Yet this behavior is not protocol-enforced—rational nodes can replace the FIFO strategy with custom rules if doing so yields financial gain. Because other nodes cannot detect or challenge such deviations, the final order remains valid under the protocol, even if manipulated.
2.3 Maximal Extractable Value (MEV)
This ability to manipulate message order gives rise to Maximal Extractable Value (MEV) [11, 12, 5], the profit a node can extract by controlling transaction sequencing. MEV is especially prevalent in decentralized finance (DeFi), where transaction order can influence token prices, arbitrage opportunities, and the outcome of smart contract interactions.
Nodes may front-run or back-run user transactions, placing their own strategically crafted messages before or after victim transactions to extract value [12]. Nodes may also observe transactions with profit potential, copy them, and submit imitated versions with higher priority [13].
Although some researchers argue that MEV can improve market efficiency [14], the dominant view is that MEV constitutes an attack vector and undermines fairness [11]. While application-level defenses exist [15, 16], they do not address the underlying vulnerability: that the message ordering mechanism is manipulable.
2.4 Permissioned and Permissionless Fair Ordering
In a permissionless setting for consensus protocols, the set of participating nodes is open and dynamic, meaning that nodes can join or leave the network frequently. In contrast, permissioned settings assume a fixed or controlled set of participants.
In the context of fair ordering protocols, there is no widely accepted definition of what it means for a fair ordering protocol to be permissionless. The closest formal notion is player replaceability [17], which captures the idea that participants in the consensus can be replaced without affecting the protocol’s security or fairness guarantees.
In practice, many fair ordering protocols designed for permissioned settings can be adapted to permissionless environments through techniques such as committee rotation or network reinitialization [18]. For example, a system can periodically select a new subset of nodes to participate in consensus or reconfigure itself when nodes join or leave. This transformation is primarily a matter of efficiency, not of fundamental design.
The main challenge in adapting fair ordering to a permissionless setting lies in the cost of reinitialization, especially when the protocol relies on threshold cryptographic schemes that require expensive bootstrap procedures such as distributed key generation. In such cases, frequent committee changes can introduce significant overhead. Nevertheless, this issue can be mitigated by reducing the frequency of committee reconfiguration, striking a balance between security and performance. Given these observations, we do not draw a strict distinction between permissioned and permissionless fair ordering protocols in this work.
3 Message Ordering Approaches and Objectives
@@
This section formalizes different message ordering rules in the Byzantine fault-tolerant setting and discusses in detail the requirements of executing these ordering rules and their limitations.
@@
We define a message ordering rule as a function that sorts a set of messages into a specific sequence based on certain metadata. Let $\mathcal{S}^{m}$ and $\mathcal{O}^{m}$ be the family of the sets of $m$ unordered messages and the lists of ordered messages ($|S|=|O|=m$), respectively. Let $\mathcal{D}$ denote the set of metadata used to order messages (varies depending on the actual ordering rule). Now we define a message ordering rule function:
@@
That is, a message ordering rule function orders a set of messages $S$, which comes from the aggregation of all local message pools of the network nodes, using some metadata $D$ and outputs the messages in an ordered list. The output $O$ is a permutation of $S$, i.e., it contains the same elements as $S$ but arranged in a specific order. Below, we discuss in detail the types of ordering rules $r$ and the corresponding metadata $D$ used to decide the ordering of the messages using these rules.
3.1 FIFO Ordering Rule
@@
However, since messages can be received by different nodes at different times and orders at the network layer, the major challenge of FIFO ordering rules is to derive a fair global message order using these different local views. Moreover, in the Byzantine fault-tolerant setting, the rules have to consider that up to $f$ local views are missing or indistinguishably malicious. The two main methods to achieve FIFO ordering under these conditions are ranked voting and median timestamping.
3.1.1 FIFO via Ranked Voting
The process of collectively agreeing on an order when each voter node has a different local order preference can be modeled by ranked-choice voting. Therefore, there are some proposals to derive the order of messages by collecting ranked local message arrival orders from each individual node and aggregating them [19, 17, 21, 22, 1, 23, 24, 37]. Different from normal ranked-choice voting aggregation, the algorithms have to account for that not all messages are received by all nodes, and therefore the local rankings are incomplete from the global perspective.
Given an unordered message set $S$, let $O^{S}_{i}$ be the ordered list of messages based on the arrival order of messages locally observed by node $i$. FIFO order via voting is defined as the following ordering rule:
@@ 3.1.1 FIFO via Ranked Voting
In FIFO via ranked voting, each node’s local order of messages acts as a “ballot” that ranks the messages, and the final ordering is decided by majority preference among nodes. However, this method may not always yield a total FIFO order due to the cases of draw Condorcet cycles [38], where cyclic preferences among nodes are found [19]. Figure 3 is a demonstration of how a Condorcet cycle is formed in ranked voting. In this example, the order of the three messages is indistinguishable because the result of the ranked voting is a draw. Even worse, such cycles can be infinitely extended to involve an arbitrary number of messages. This raises an attack opportunity of deliberately creating or extending Condorcet cycles in the case where malicious external message senders submit messages to different nodes at carefully designed times [25].
@@
Figure 3: An example of a Condorcet cycle from ranked voting. Among the three nodes, two nodes receive message 1 before message 2, two nodes receive message 2 before message 3, also two nodes receive message 3 before message 1.
@@
To tolerate Condorcet cycles, different relaxed notions of order fairness are proposed. Block order fairness [19] (later referred to as batch order fairness) and similarly differential order fairness [21] are proposed that allow multiple messages to be deemed as received at the same time. In this case, any ordering among these messages is deemed acceptable. For completeness, the definition is provided below:
Definition 1 ($\gamma$-Batch Order Fairness)
Let $m_{1}$ and $m_{2}$ be two messages that are both received and processed by all honest nodes. Let each honest node maintain a local reception order over messages. We say that a protocol satisfies $\gamma$-batch order fairness for some $\gamma\in(0.5,1]$ if the following holds: If at least a $\gamma$-fraction of honest nodes observed $m_{1}$ before $m_{2}$ in their local reception order, then all honest nodes output $m_{1}$ no later than $m_{2}$.
However, this definition is under the setting that messages can be output in batch, in which messages are regarded as indistinguishable in their order. This does not apply to the setting of blockchain, where a total ordering must be derived. Practically, this means block proposers still have the freedom to order messages in batches to maximize personal utilities.
@@
Moreover, as Condorcet cycles can be arbitrarily long, making sure messages in Condorcet cycles commit at the same time also impacts the liveness of the protocol [19]. A solution to this problem is to assume an external synchrony time bound [19], so that when the time of receiving two messages is over the time bound, they must not be in a Condorcet cycle. The drawback of this method is the high latency to commit a message and the requirement of an additional external synchrony assumption. To improve on that, one way is to use ranked pairs [39] to break Condorcet cycles without violating FIFO order fairness [25, 26]. Themis [1] also introduces another method called batch-unspooling that commits messages that are part of a cycle in a round without requiring observing the full cycle, and allows remaining messages in the cycle to commit in following rounds.
@@
To summarize, this line of research attempts to handle Condorcet cycles in a way that minimizes their effect on latency. However, taking a step back, improvements can also be made by actively preventing the formation of Condorcet cycles. Therefore, a potential research direction that can advance FIFO via voting proposals is to reduce the probability of forming Condorcet cycles, for example, by requiring more frequent message propagation [25].
3.1.2 FIFO via Timestamping